# The Efficacy of Semantics-Preserving Transformations in Self-Supervised Learning for Medical Ultrasound

### v1

@@ Appendix A Dataset Details
This section provides further details regarding the composition of the LUSData and COVIDx-US datasets, stratified by different attributes. Table XII provides a breakdown of the characteristics of the unlabelled data, training set, validation set, local test set, and external test set. As can be seen in the table, the external test set’s manufacturer and probe type distributions differ greatly from those in the local test set. As such, the external test set constitutes a meaningful assessment of generalizability for models trained using the LUSData dataset. Table XIII provides similar information for the COVIDx-US dataset [23]. Note that the probe types listed in Tables XII and XIII are predictions produced by a ultrasound cleaning software (UltraMask, Deep Breathe Inc., London, ON, Canada) – not from meta-data accompanying the examinations.
TABLE XII: Characteristics of the ultrasound videos in the LUSData dataset. The number of videos possessing known values for a variety of attributes are displayed. Percentages of the total in each split are provided as well, but some do not sum to $100$ due to rounding.
@@
TABLE XIII: Characteristics of the ultrasound videos in the COVIDx-US dataset. The number of videos possessing known values for a variety of attributes are displayed. Percentages of the total in each split are provided as well, but some do not sum to $100$ due to rounding.
@@ Appendix B Transformation Runtime Estimates
Appendix C BYOL Transformations
We investigated the data augmentation pipeline from the BYOL study [13]. More specifically, we adopted the symmetric version from the VICReg paper [39], which uses the same transformations and probabilities for each of the two branches in the joint embedding architecture. The transformations and their parameters are widely adopted; nevertheless, we detail their operation and parameters below.
C-A Crop and Resize (B00)
A rectangular crop of the input image is designated at a random location. The area of the cropped region is sampled from the uniform distribution $\mathcal{U}(0.08,1)$. The cropped region’s aspect ratio is sampled from the uniform distribution $\mathcal{U}(0.75,1.\overline{33})$,. Its width and height are calculated accordingly. The cropped region is then resized to the original image dimension.
C-B Horizontal Reflection (B01)
@@ C-C Color Jitter (B02)
C-D Conversion to Grayscale (B03)
Images are converted to grayscale. The output images has three channels, such that each channel has the same pixel intensity.
C-E Gaussian Blur (B04)
@@ Appendix D Ultrasound-Specific Transformations
Several of the transformations operate on the pixels contained within the ultrasound (US) beam. As such, the geometrical form of the beam was required to perform some transformations. We adopted the same naming convention for the vertices of the US beam as Kim et al. [40]. Let $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ represent the respective locations of the top left, top right, bottom left, and bottom right vertices, and let $\langle x_{i},y_{i}\rangle$ be the x- and y-coordinates of $p_{i}$ in image space. For convex beam shapes, we denote the intersection of lines $\overleftrightarrow{p_{1}p_{3}}$ and $\overleftrightarrow{p_{2}p_{4}}$ as $p_{0}$. Fig. 9 depicts the arrangement of these points for each of the three main US beam shapes: linear, curvilinear, and phased array. A software tool was used to estimate the beam vertices and probe type for all videos in each dataset (UltraMask, Deep Breathe Inc., Canada).
D-A Probe Type Change (U00)
To produce a transformed ultrasound image with a different beam shape, a mapping that gives the location of pixels in the original image for each coordinate in the new image is calculated. Concretely, the function $\mathbf{f}:\mathbb{R}^{2}\rightarrow[-1,1]^{2}$ returns the coordinates of the point in the original image that corresponds to a point in the transformed image. Note that $(-1,-1)$ corresponds to the top left of the original image. Pixel intensities in the transformed image are interpolated according to their corresponding location in the original image.
Figure 9: Location of the named beam vertices for each of the three main beam shapes in US imaging
Algorithm 1 details the calculation of $\mathbf{f}_{\ell\rightarrow c}$ for converting linear beams to convex beams with a random radius factor $\rho\sim\mathcal{U}(1,2)$, along with new beam vertices. Similarly, curvilinear and phased array beam shapes are converted to linear beam shapes. Algorithm 2 details the calculation of the mapping $\mathbf{f}_{c\rightarrow\ell}$ that transforms convex beam shapes to linear shapes, along with calculations for the updated named coordinates. To ensure that no aspects of the old beam remain on the image, bitmask $M^{\prime}\in\{0,1\}^{h\times w}$ is produced using the new named coordinates.
Since the private dataset was resized to square images that exactly encapsulated the beam, images are resized to match their original aspect ratios to ensure that the sectors were circular. They are then resized to their original dimensions following the transformation.
Algorithm 1 Compute a point mapping for linear to curvilinear beam shape, along with new beam vertices
0:  Beam vertices $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$; radius factor $\rho$; coordinate maps $\mathbf{x}=\mathbf{1}_{h\times 1}[0,1,\ldots,w-1]$ and         $\mathbf{y}=[0,1,\ldots,h-1]^{T}\mathbf{1}_{1\times w}$
@@ D-A Probe Type Change (U00)
7:  $x_{2}^{\prime}\leftarrow 2x_{0}^{\prime}-x_{1}^{\prime}$
@@
10:  $\mathbf{f}_{\ell\rightarrow c}\leftarrow\begin{pmatrix}\frac{\bm{\phi}+(x_{i}-w/2)/w}{|\,\bm{\phi}[y_{3}^{\prime},0]\,|}\\[7.0pt] \frac{2\sqrt{(x_{0}^{\prime}-\mathbf{x})^{2}+(y_{0}^{\prime}-\mathbf{y})^{2}}-r_{t}}{r_{b}-r_{t}}-1\end{pmatrix}$
11:  return  $\mathbf{f}_{\ell\rightarrow c}$, $p_{1}^{\prime}$, $p_{2}^{\prime}$, $p_{3}^{\prime}$, $p_{4}^{\prime}$
Algorithm 2 Compute a point mapping for convex to linear beam shape, along with new beam vertices
0:  Beam vertices $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$; point of intersection $p_{0}$; angle of intersection $\theta_{0}$; width fraction $\omega\in[0,1]$; coordinate maps $\mathbf{x}=\mathbf{1}_{h\times 1}[0,1,\ldots,w-1]$ and         $\mathbf{y}=[0,1,\ldots,h-1]^{T}\mathbf{1}_{1\times w}$
@@
14:  return  $\mathbf{f}_{c\rightarrow\ell}$, $p_{1}^{\prime}$, $p_{2}^{\prime}$, $p_{3}^{\prime}$, $p_{4}^{\prime}$
D-B Convexity Change (U01)
To mimic an alternative convex beam shape with a different $\theta_{0}$, a mapping is calculated that results in a new beam shape wherein $p_{0}$ is translated vertically. A new value for the width of the top of the beam is randomly calculated, facilitating the specification of a new $p_{0}$. Given the new $p_{0}$, a pixel map $\mathbf{f}_{c\rightarrow c^{\prime}}$ is computed according to Algorithm 3. Similar to the probe type change transformation, pixel intensities at each coordinate in the transformed image are interpolated according to the corresponding coordinate in the original image returned by $\mathbf{f}_{c\rightarrow c^{\prime}}$.
Algorithm 3 Compute a point mapping from an original to a modified convex beam shape.
0:  Beam vertices $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$; point of intersection $p_{0}$; angle of intersection $\theta_{0}$; new top width $w^{\prime}$; coordinate maps $\mathbf{x}=\mathbf{1}_{h\times 1}[0,1,\ldots,w-1]$ and         $\mathbf{y}=[0,1,\ldots,h-1]^{T}\mathbf{1}_{1\times w}$
@@ D-B Convexity Change (U01)
15:  return  $\mathbf{f}_{c\rightarrow c^{\prime}}$, $p_{1}^{\prime}$, $p_{2}^{\prime}$, $p_{3}^{\prime}$, $p_{4}^{\prime}$
D-C Wavelet Denoising (U02)
@@
D-D Contrast-Limited Adaptive Histogram Equalization (U03)
@@ D-E Gamma Correction (U04)
D-F Brightness and Contrast Change (U05)
The brightness and contrast of the image are modified. The brightness change factor, contrast change factor are sampled from $\mathcal{U}(0.6,1.4)$, $\mathcal{U}(0.6,1.4)$, respectively. The image is then multiplied element-wise by its beam mask, to ensure black regions external to the beam remain black.
D-G Depth Change Simulation (U06)
preserves the centre for linear beam shapes and preserves $p_{0}$ for convex beam shapes. The magnitude of the zoom transformation, $d$, is randomly sampled from a uniform distribution. Increasing the depth corresponds to zooming out ($d>1$), while decreasing the depth corresponds to zooming in ($d<1$).
D-H Speckle Noise Simulation (U07)
Following Singh et al.’s method [28], this transformation calculates synthetic speckle noise and applies it to the ultrasound beam. Various parameters of the algorithm are randomly sampled upon each invocation. The lateral and axial resolutions for interpolation are random integers drawn from the ranges $[35,45]$ and $[75,85]$, respectively. The number of synthetic phasors is randomly drawn from the integer range $[5,10]$. Sample points on the image are evenly spaced in Cartesian coordinates for linear beam shapes. For convex beams, the sample points are evenly spaced in polar coordinates.
D-I Gaussian Noise Simulation (U08)
@@
D-J Salt and Pepper Noise Simulation (U09)
A random assortment of points in the image are set to to $255$ (salt) or 00 (pepper). The fractions of pixels set to salt and pepper values are sampled randomly according to $\mathcal{U}(0.001,0.005)$.
D-K Horizontal Reflection (U10)
@@ Appendix E Pleural Line Object Detection Training
The Single Shot Detector (SSD) method [34] was employed to perform a cursory evaluation of the pretrained models on an object detection task. The PL task involved localizing instances of the pleural line in lung US images, which can be described as a bright horizontal line that is typically situated slightly below the vertical level of the ribs. It is only visible between the rib spaces, since bone blocks the ultrasound scan lines. The artifact represents the interface between the parietal and viesceral linings of the lung.
As in the classification experiments, we used the MobileNetV3Small architecture as the backbone of the network. The feature maps outputted from a designated set of layers were passed to the SSD regression and classification heads. We selected a range of layers whose feature maps had varying spatial resolution. Table XIV provides the identities and dimensions of the feature maps from the backbone that were fed to the SSD model heads. The set of default anchor box aspect ratios was manually specified after examining the distribution of bounding box aspect ratios in the training set. The 25th percentile was $2.894$, and the 75th percentile was $4.989$. The pleural line artifact typically has a much greater width than height. Accordingly, we designated the set of default anchor box aspect ratios ($w/h$) as $\{1,2,3,4,5\}$. Six anchor box scales were used. The first five were spaced out evenly over the range $[0.023,0.170]$, which correspond to the square roots of the minimum and maximum areas of the bounding box labels present in the training set, in normalized image coordinates. The final scale is $1$, which is included by default.
@@
TABLE XIV: MobileNetV3Small block indices and the corresponding dimensions of the feature maps that they output, given an input of size $128\times 128\times 3$.
@@ Appendix F Leave-one-out Analysis Statistical Testing
As outlined in Section IV-A, statistical testing was performed to detect differences between pretrained models trained using an ablated version of the BYOL and AugUS-O pipelines and baseline models pretrained on the original pipelines. Each ablated version of the pipeline was missing one transformation from the data augmentation pipeline. Ten-fold cross-validation conducted on the training set provided $10$ samples of test AUC metrics for both the A-line versus B-line (AB) and pleural effusion (PE) binary classification tasks. The samples were taken as a proxy for test-time performance for linear classifiers trained on each of the above downstream tasks.
To determine whether the mean test AUC for each ablated model was different from the baseline model, hypothesis testing was conducted. The model pretrained using the original pipeline was the control group, while the models pretrained using ablated vesions of the pipeline were the experimental groups. First, Friedman’s test [35] was conducted to determine if there was any difference in the mean test AUC among the baseline and ablated models. We selected a nonparametric multiple comparisons test because of the lack of assumptions regarding normality or homogeneity of variances. Each collection had $10$ samples. Table XV details the results of Friedman’s test for each pipeline and classification task. Friedman’s test detected differences among the collection of test AUC for both classification tasks with the BYOL pipeline. Only the AB task exhibited significant differences for the AugUS-O pipeline.
TABLE XV: Friedman test statistics and p-values for mean cross-validation test AUC attained by models pretrained using an entire data augmentation pipeline and ablated versions of it.
@@
When the null hypothesis of the Friedman test was rejected, post-hoc tests were conducted to determine whether any of the test AUC means in the experimental groups were significantly different than the control group. The Wilcoxon Sign-Rank Test [36] was designated as the post-hoc test, due to its absence of any normality assumptions. Note that for each pipeline, $n$ comparisons were performed, where $n$ is the number of transformations within the pipeline. The Holm-Bonferroni correction [37] was applied to keep the family-wise error rate at $\alpha=0.05$ for each pipeline/task combination. Results of the post-hoc tests are given in Table XVI. No post-hoc tests were performed for the AugUS-O pipeline evaluated on the PE task because the Friedman test revealed no significant differences.
TABLE XVI: Test statistics ($T$) and p-values obtained from the Wilcoxon Sign-Rank post-hoc tests that compared linear classifiers trained with ablated models’ features to a control linear classifier trained on the baseline model. Experimental groups are identified according to the left-out transformation, as defined in Tables II and III.

### v2

@@ 5 Conclusion
CRediT authorship contribution statement
Blake VanBerlo: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project Administration, Funding Acquisition Jesse Hoey: Conceptualization, Writing - Review & Editing, Supervision Alexander Wong: Conceptualization, Writing - Review & Editing, Supervision Robert Arntfield: Resources, Data Curation, Writing - Review & Editing
Declaration of competing interest
There are no competing interests for all authors.
Data availability
The LUSData dataset is private. COVIDxUS [26] is publicly available. Code will be published in a public GitHub repository upon publication.
Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council of Canada, as Blake VanBerlo is a Vanier Scholar (FRN 186945). Computational resource support was also provided by Compute Ontario (computeontario.ca) and the Digital Research Alliance of Canada
(alliance.can.ca).
@@ Appendix A Dataset Details
This section provides further details regarding the composition of the LUSData and COVIDx-US datasets, stratified by different attributes. Table 10 provides a breakdown of the characteristics of the unlabelled data, training set, validation set, local test set, and external test set. As can be seen in the table, the external test set’s manufacturer and probe type distributions differ greatly from those in the local test set. As such, the external test set constitutes a meaningful assessment of generalizability for models trained using the LUSData dataset. Table 11 provides similar information for the COVIDx-US dataset [26]. Note that the probe types listed in Tables 10 and 11 are predictions produced by a product and not from meta-data accompanying the examinations. [^5] Product name withheld to preserve anonymity during peer review.
Table 10: Characteristics of the ultrasound videos in the LUSData dataset. The number of videos possessing known values for a variety of attributes are displayed. Percentages of the total in each split are provided as well, but some do not sum to $100$ due to rounding.
@@
Table 11: Characteristics of the ultrasound videos in the COVIDx-US dataset. The number of videos possessing known values for a variety of attributes are displayed. Percentages of the total in each split are provided as well, but some do not sum to $100$ due to rounding.
@@ Appendix B Transformation Runtime Estimates
Appendix C StandardAug Transformations
We investigated a standard data augmentation pipeline that has used extensively in the SSL literature [13, 14, 15, 16]. To standardize experiments, we adopted the symmetric version from the VICReg paper [16], which uses the same transformations and probabilities for each of the two branches in the joint embedding architecture. The transformations and their parameter settings are widely adopted. Below, we detail their operation and parameter assignments.
C.1 Crop and Resize (B00)
A rectangular crop of the input image is designated at a random location. The area of the cropped region is sampled from the uniform distribution $\mathcal{U}(0.08,1)$. The cropped region’s aspect ratio is sampled from the uniform distribution $\mathcal{U}(0.75,1.33)$. Its width and height are calculated accordingly. The cropped region is then resized to the original image dimension.
C.2 Horizontal Reflection (B01)
@@ C.3 Color Jitter (B02)
C.4 Conversion to Grayscale (B03)
Images are converted to grayscale. The output images have three channels, such that each channel has the same pixel intensity.
C.5 Gaussian Blur (B04)
@@ Appendix D Ultrasound-Specific Transformations
Several of the transformations operate on the pixels contained within the ultrasound field of view (FOV). As such, the geometrical form of the FOV was required to perform some transformations. We adopted the same naming convention for the vertices of the ultrasound FOV as Kim et al. [42]. Let $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ represent the respective locations of the top left, top right, bottom left, and bottom right vertices, and let $\langle x_{i},y_{i}\rangle$ be the x- and y-coordinates of $p_{i}$ in image space. For convex FOV shapes, we denote the intersection of lines $\overleftrightarrow{p_{1}p_{3}}$ and $\overleftrightarrow{p_{2}p_{4}}$ as $p_{0}$. Fig. 9 depicts the arrangement of these points for each of the three main ultrasound FOV shapes: linear, curvilinear, and phased array. A software tool was used to estimate the FOV shape and probe type for all videos in each dataset (UltraMask, Deep Breathe Inc., Canada).
D.1 Probe Type Change (U00)
To produce a transformed ultrasound image with a different FOV shape, a mapping that gives the location of pixels in the original image for each coordinate in the new image is calculated. Concretely, the function $\mathbf{f}:\mathbb{R}^{2}\rightarrow[-1,1]^{2}$ returns the coordinates of the point in the original image that corresponds to a point in the transformed image. Note that $(-1,-1)$ corresponds to the top left of the original image. Pixel intensities in the transformed image are interpolated according to their corresponding location in the original image.
Figure 9: Locations of the named FOV vertices for each of the three main field of view shapes in US imaging
Algorithm 1 details the calculation of $\mathbf{f}_{\ell\rightarrow c}$ for converting linear FOVs to convex FOVs with a random radius factor $\rho\sim\mathcal{U}(1,2)$, along with new FOV vertices. Similarly, curvilinear and phased array FOV shapes are converted to linear FOV shapes. Algorithm 2 details the calculation of the mapping $\mathbf{f}_{c\rightarrow\ell}$ that transforms convex FOV shapes to linear shapes, along with calculations for the updated named coordinates. To ensure that no aspects of the old FOV remain on the image, bitmask $M^{\prime}\in\{0,1\}^{h\times w}$ is produced using the new named coordinates.
Since the private dataset was resized to square images that exactly encapsulated the FOV, images are resized to match their original aspect ratios to ensure that the sectors were circular. They are then resized to their original dimensions following the transformation.
Algorithm 1 Compute a point mapping for linear to curvilinear FOV shape, along with new FOV vertices
0:  FOV vertices $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$; radius factor $\rho$; coordinate maps $\mathbf{x}=\mathbf{1}_{h\times 1}[0,1,\ldots,w-1]$ and $\mathbf{y}=[0,1,\ldots,h-1]^{T}\mathbf{1}_{1\times w}$
@@ D.1 Probe Type Change (U00)
7:  $x_{2}^{\prime}\leftarrow x_{0}^{\prime}+(y_{1}-y_{0}^{\prime})(x_{0}^{\prime}-x_{3})/(y_{3}^{\prime}-y_{0}^{\prime})$
@@
[…]
Algorithm 2 Compute a point mapping for convex to linear FOV shape, along with new FOV vertices
0:  FOV vertices $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$; point of intersection $p_{0}$; angle of intersection $\theta_{0}$; width fraction $\omega\in[0,1]$; coordinate maps $\mathbf{x}=\mathbf{1}_{h\times 1}[0,1,\ldots,w-1]$ and $\mathbf{y}=[0,1,\ldots,h-1]^{T}\mathbf{1}_{1\times w}$
@@
[…]
D.2 Convexity Change (U01)
To mimic an alternative convex FOV shape with a different $\theta_{0}$, a mapping is calculated that results in a new FOV shape wherein $p_{0}$ is translated vertically. A new value for the width of the top of the FOV is randomly calculated, facilitating the specification of a new $p_{0}$. Given the new $p_{0}$, a pixel map $\mathbf{f}_{c\rightarrow c^{\prime}}$ is computed according to Algorithm 3. Similar to the probe type change transformation, pixel intensities at each coordinate in the transformed image are interpolated according to the corresponding coordinate in the original image returned by $\mathbf{f}_{c\rightarrow c^{\prime}}$.
Algorithm 3 Compute a point mapping from an original to a modified convex FOV shape.
0:  FOV vertices $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$; point of intersection $p_{0}$; angle of intersection $\theta_{0}$; new top width $w^{\prime}$; coordinate maps $\mathbf{x}=\mathbf{1}_{h\times 1}[0,1,\ldots,w-1]$ and $\mathbf{y}=[0,1,\ldots,h-1]^{T}\mathbf{1}_{1\times w}$
@@ D.2 Convexity Change (U01)
[…]
D.3 Wavelet Denoising (U02)
@@
D.4 Contrast-Limited Adaptive Histogram Equalization (U03)
@@ D.5 Gamma Correction (U04)
D.6 Brightness and Contrast Change (U05)
The brightness and contrast of the image are modified. The brightness change factor, contrast change factor are sampled from $\mathcal{U}(0.6,1.4)$, $\mathcal{U}(0.6,1.4)$, respectively. The image is then multiplied element-wise by its FOV mask, to ensure black regions external to the FOV remain black.
D.7 Depth Change Simulation (U06)
preserves the centre for linear FOV shapes and preserves $p_{0}$ for convex FOV shapes. The magnitude of the zoom transformation, $d$, is randomly sampled from a uniform distribution. Increasing the depth corresponds to zooming out ($d>1$), while decreasing the depth corresponds to zooming in ($d<1$).
D.8 Speckle Noise Simulation (U07)
Following Singh et al.’s method [31], this transformation calculates synthetic speckle noise and applies it to the ultrasound FOV. Various parameters of the algorithm are randomly sampled upon each invocation. The lateral and axial resolutions for interpolation are random integers drawn from the ranges $[35,45]$ and $[75,85]$, respectively. The number of synthetic phasors is randomly drawn from the integer range $[5,10]$. Sample points on the image are evenly spaced in Cartesian coordinates for linear FOV shapes. For convex FOVs, the sample points are evenly spaced in polar coordinates.
D.9 Gaussian Noise Simulation (U08)
@@
D.10 Salt and Pepper Noise Simulation (U09)
A random assortment of points in the image are set to $255$ (salt) or 00 (pepper). The fractions of pixels set to salt and pepper values are sampled randomly according to $\mathcal{U}(0.001,0.005)$.
D.11 Horizontal Reflection (U10)
@@ Appendix E Pleural Line Object Detection Training
The Single Shot Detector (SSD) method [36] was employed to perform a cursory evaluation of the pretrained models on an object detection task. The PL task involved localizing instances of the pleural line in lung US images, which can be described as a bright horizontal line that is typically situated slightly below the vertical level of the ribs. It is only visible between the rib spaces, since bone blocks the ultrasound scan lines. The artifact represents the interface between the parietal and visceral linings of the lung. In a properly acquired lung US view,
As in the classification experiments, we used the MobileNetV3Small architecture as the backbone of the network. There is precedent for using the SSD object detection method, as it has been applied to assess the object detection capabilities of MobileNet architectures [43, 32]. The feature maps outputted from a designated set of layers were passed to the SSD regression and classification heads. We selected a range of layers whose feature maps had varying spatial resolution. Table 12 provides the identities and dimensions of the feature maps from the backbone that were fed to the SSD model head. The head contained 43 7964379643\,79643 796 trainable parameters, which was light compared to the backbone (927 008927008927\,008927 008).
The set of default anchor box aspect ratios was manually specified after examining the distribution of bounding box aspect ratios in the training set. The 25th percentile was $2.894$, and the 75th percentile was $4.989$. The pleural line typically has a much greater width than height. Accordingly, we designated the set of default anchor box aspect ratios ($w/h$) as $\{1,2,3,4,5\}$. Six anchor box scales were used. The first five were spaced out evenly over the range $[0.023,0.170]$, which correspond to the square roots of the minimum and maximum areas of the bounding box labels present in the training set, in normalized image coordinates. The final scale is $1.000$, which is included by default. The box confidence threshold was $0.01$ and the intersection over union threshold to match anchors to box labels was $0.3$. The non-maximum suppression (NMS) threshold was $0.45$ and the number of detected boxes to keep after NMS was $50$.
@@
Table 12: MobileNetV3Small block indices and the corresponding dimensions of the feature maps that they output, given an input of size $128\times 128\times 3$.
@@ Appendix F Leave-one-out Analysis Statistical Testing
As outlined in Section IV-A, statistical testing was performed to detect differences between pretrained models trained using an ablated version of the StandardAug and AugUS-O pipelines and baseline models pretrained on the original pipelines. Each ablated version of the pipeline was missing one transformation from the data augmentation pipeline. Ten-fold cross-validation conducted on the training set provided $10$ samples of test AUC metrics for both the A-line versus B-line (AB) and pleural effusion (PE) binary classification tasks. The samples were taken as a proxy for test-time performance for linear classifiers trained on each of the above downstream tasks.
To determine whether the mean test AUC for each ablated model was different from the baseline model, hypothesis testing was conducted. The model pretrained using the original pipeline was the control group, while the models pretrained using ablated versions of the pipeline were the experimental groups. First, Friedman’s test [37] was conducted to determine if there was any difference in the mean test AUC among the baseline and ablated models. We selected a nonparametric multiple comparisons test because of the lack of assumptions regarding normality or homogeneity of variances. Each collection had $10$ samples. Table 13 details the results of Friedman’s test for each pipeline and classification task. Friedman’s test detected differences among the collection of test AUC for both classification tasks with the StandardAug pipeline. Only the AB task exhibited significant differences for the AugUS-O pipeline.
Table 13: Friedman test statistics and p-values for mean cross-validation test AUC attained by models pretrained using an entire data augmentation pipeline and ablated versions of it.
@@
When the null hypothesis of the Friedman test was rejected, post-hoc tests were conducted to determine whether any of the test AUC means in the experimental groups were significantly different than the control group. The Wilcoxon Sign-Rank Test [38] was designated as the post-hoc test, due to its absence of any normality assumptions. Note that for each pipeline, $n$ comparisons were performed, where $n$ is the number of transformations within the pipeline. The Holm-Bonferroni correction [39] was applied to keep the family-wise error rate at $\alpha=0.05$ for each pipeline/task combination. Results of the post-hoc tests are given in Table 14. No post-hoc tests were performed for the AugUS-O pipeline evaluated on the PE task because the Friedman test revealed no significant differences.
Table 14: Test statistics ($T$) and p-values obtained from the Wilcoxon Sign-Rank post-hoc tests that compared linear classifiers trained with ablated models’ features to a control linear classifier trained on the baseline model. Experimental groups are identified according to the left-out transformation, as defined in Tables II and III.